OpenStudy (anonymous):

A swimming pool is initially shocked with chlorine to bring the chlorine concentration to 5 ppm (parts per million). Chlorine dissipates in reaction to bacteria and sun at a rate of about 15% per day. Above a chlorine concentration of 2 ppm, swimmers experience burning eyes, and below a concentration of 1 ppm, bacteria and algae start to proliferate in the pool environment. (a) Construct an exponential decay function that describes the chlorine concentration (in parts per million) over time.

OpenStudy (mathstudent55):

Exponential decay is given by y = x(1 - r)^t where x is the starting amount (concentration) r is the rate of decay (written as a decimal) t = time

OpenStudy (kainui):

Is this a calculus or chemistry class? I can make this very in depth or not very depending.

OpenStudy (anonymous):

it is college algebra actually.

OpenStudy (anonymous):

should it be y=5(0.85)^x?

OpenStudy (perl):

y = 5 ( 1 - .15) ^t

OpenStudy (anonymous):

This makes sense..theres more to the question... c. How many days will it take for the chlorine to reach a level tolerable for swimmers? How many days before bacteria and algea will start to grow and you will need to add more chlorine? Round your answers to one decimal place. It will take __________days for the chlorine to reach a level tolerable for swimmers. It will take ________ days before bacteria and algea start to grow.

OpenStudy (anonymous):

i apologize for the inconvenience

OpenStudy (perl):

what was part b)

OpenStudy (anonymous):

part be was simply to complete construct a table based on the equation we already formulated...

OpenStudy (perl):

ok thanks, give a sec

OpenStudy (perl):

so you want to know when is y less than 2

OpenStudy (anonymous):

yes...it says to round to one decimal...

OpenStudy (anonymous):

oh so it should be (6 and 1.8)

OpenStudy (perl):

you know, lets solve by logs

OpenStudy (anonymous):

can you help me on logs...hehe im not really good at that..

OpenStudy (perl):

2 = 5 (.85)^t 2/5 = (.85)^t now take logs of both sides

OpenStudy (anonymous):

i got t=5.6

OpenStudy (perl):

ln (2/5) = ln ( .85^t) ln (2/5) = t ln (.85)

OpenStudy (perl):

yes I got that

OpenStudy (anonymous):

is t=5.6 right?

OpenStudy (perl):

so after t > 5.638 , the pool is tolerable for humans

OpenStudy (perl):

5.368 days

OpenStudy (anonymous):

oku mean 5.638 or 5.368?

OpenStudy (perl):

5.638, woops

OpenStudy (anonymous):

what about for this part of the question? It will take ________ days before bacteria and algea start to grow.

OpenStudy (perl):

so by 6 days it is tolerable , after shocking the pool

OpenStudy (perl):

now we have to solve , when is y > 1 , so we can solve 1 = 5 (.85 ^t)

OpenStudy (anonymous):

t=9.9 right :)

OpenStudy (anonymous):

thank you so much i got them right...im using wiley plus and it is difficult for me...this website is great help..

OpenStudy (perl):

:D

OpenStudy (anonymous):

im done with homework thanks to your help, see you another time :) have a great week..

OpenStudy (perl):

goodbye